# Repeating Flares, X-ray Outbursts and Delayed Infrared Emission: A Comprehensive Compilation of Optical Tidal Disruption Events

### v1

@@ 3.3 Ultraviolet Photometry
To evaluate TDE variability in UV-optical bands, we used data from the Swift Ultraviolet and Optical Telescope (UVOT) (Roming et al. 2005), which provides photometry in three near-ultraviolet (UVW2, UVM2 and UVW1) and three optical (U, B, V) bands.
UVOT data were downloaded from HEASARC101010https://heasarc.gsfc.nasa.gov/ and reduced using standard procedures111111http://www.swift.ac.uk/analysis/uvot/image.php (HEAsoft package v. 6.33.2). After checking the correct World Coordinate System alignment with USNO-B Catalogue (Monet et al. 2003), we combined image extensions using uvotimsum and merged exposure maps.
Sources were detected using uvotdetect. As with XRT data (see Appendix B.1), we assigned UVOT counterparts based on proximity to the TDE source coordinates: if a UVOT sources was detected within , its coordinates were adopted. If no source was detected within , we used the TDE sources coordinates.
We performed source photometry using uvotsource with an aperture radius of for all filters. The background region was an annulus centred at the source with inner and outer radii of and , respectively, removing emission from overlapping sources. All magnitudes were corrected for Galactic extinction using reddening estimates from Schlafly & Finkbeiner (2011) and the extinction model from Fitzpatrick & Massa (2007). We note that these data are not host subtracted.
@@ 3.4 Spectroscopic data
All optical spectra included in the catalogue were obtained from either TNS or the Weizmann Interactive Supernova Data Repository121212https://www.wiserep.org (WISeREP; Yaron & Gal-Yam 2012). Hence, the catalogue does not include optical spectra for TDEs not classified in TNS. For the 12 TDEs from Hammerstein et al. (2023), the classification spectra can be found in their Figs. 2 and 14-16. For AT 2018lni, AT 2019cho, AT 2019mha and AT 2019meg, the classification spectra can be found in van Velzen et al. (2021).
Additionally, we list references for (mainly optical) spectra of 13 TDEs that are not in the TNS-TDE sample or included in a previously published catalogue.
AT 2023vto: Fig. 4 in Kumar et al. (2024)
AT 2022agi: Figs. 2 (optical), 3 (UV) in Sun et al. (2024)
AT 2017eqx: Figs. 1, 3 in Nicholl et al. (2019)
LSQ12dyw†131313† signifies the sources that have spectra available in WISeREP.
PTF09djl†: Figs. 4, 14 in Arcavi et al. (2014)
PTF09ge†: Fig. 12, 14 in Arcavi et al. (2014)
PS1-10jh†: Fig. 1 in Gezari et al. (2012)
iPTF15af†: Figs. 3 (optical), 4 (UV) in Blagorodnova et al. (2019)
iPTF16axa†: Figs. 7, 8 in Hung et al. (2017)
ASASSN-14li†: Fig. 3 in Holoien et al. (2016b)
ASASSN-14ae†: Figs. 4, 5 in Holoien et al. (2014)
ASASSN-15oi†: Fig. 3 in Holoien et al. (2016a)
@@ 4 Exploring the catalogue
After having constructed our main sample, we now explore its various subcategories and interesting objects. Specifically, we present the statistics and spectral classes of the TDEs in the catalogue, along with new results on TDEs with repeating flares, delayed infrared emission and X-ray outbursts.
@@ 4.1 Catalogue statistics
In this section we analyse the flaring characteristics of the TDE population, in terms of rise/decay times and the event durations. To achieve this, we apply the Bayesian Blocks (Scargle 1998; Scargle et al. 2013) algorithm which partitions a one-dimensional flux time-series into blocks, each modelled by a constant flux. To avoid over-fitting, we include a user-specified penalty parameter for each additional block.
Given flux measurements and their uncertainties , we first compute cumulative weighted sums (with ) up to each index: . These cumulative sums allow efficient computation of sums over any sub-interval via . Consequently, the maximum-likelihood estimate for the constant flux in a block spanning is .
@@
Moreover, the algorithm uses dynamic programming to determine the optimal segmentation. For each index (ranging from 1 to the total number of data points ), all possible previous boundaries (from to ) are considered. We compute:
@@
penalty is a user-specified constant subtracted each time a new block is added. We chose a default value of 10 for the penalty to avoid over-fitting or under-fitting (i.e. small or large penalty values respectively).
@@
The optimal segmentation is obtained by choosing the that maximizes candidate_score followed by a backtracking step (from to ) to recover the optimal boundaries.
@@
Applying the Bayesian Blocks algorithm to the flux light curves yields blocks within which the flux is assumed to be constant. For each block, we record its mean flux value. The peak of the flare is then identified as the block with the highest mean flux. We primarily use ZTF(zg) light curves (76/103). In case ZTF data are not available or the ZTF(zg) light curve is poorly sampled, we use the light curve with the best sampling. This results in 11 ZTF(zr) light curves, 6 ATLAS light curves, 5 ASAS-SN light curves, 1 GSA(G) light curve, 1 PS1 light curve, 1 CRTS(CL) light curve and 2 PTF light curves.
@@
To estimate the rise and decay times of the flare, we apply a threshold relative to the peak flux. Starting from the peak block, we move to lower flux levels (to lower or higher block indices) until encountering a block with a mean flux below this threshold. The first sub-threshold block on the left marks the start of the rise, while the first sub-threshold segment on the right marks the end of the decay. From these boundaries, we calculate the rise time, , and decay time, , and determine their ratio .
@@
An example of this process is displayed in Fig. 1. For this example, the rise time is 25698 days, the decay time 582192 days and the peak time 5980322 MJD. The duration of the flare, calculated as the sum of the rise and decay times, is roughly 838 days.
@@
We applied this method to 103 TDE flares in our sample, excluding most of the sources detected in 2024 and some older ones where rise or decay times could not be reliably estimated due to large gaps in the light curves. We also included both flares for AT 2022dbl and AT 2020vdq (see Sect. 4.3). Figure 2 shows the distribution of the common logarithm (i.e. logarithm with base 10) of the ratio in our sample. A normal distribution fit to this data yields and mean of -0.49, with a Kolmogorov-Smirnov (K-S) test p-value of 0.94 (K-S statistic = 0.05). This means that the ratio distribution is consistent with a log-normal distribution.
@@
Figure 3 displays the duration, rise and decay times of the flares. Gaussian fits to these distributions yield means of 35020 days, 845 days and 25020 days for the duration, and (the error is the standard error of the mean) and standard deviations of 207 days, 55 day and 167 days respectively. The p-values are 0.91, 0.85 and 0.70 (K-S statistic of 0.054, 0.059 and 0.068) for the durations, and respectively, consistent with log-normal distributions.
@@
The bottom right panel of Fig. 3 shows a scatter plot of with error bars calculated as , where is half the width of the rise/decay blocks and is half the width of the peak block. Given errors in both the independent () and dependent () variables, we employ the orthogonal distance regression (ODR) to fit a best-fit line while accounting for both uncertainties. The best-fit ODR line is shown as the red solid line in the bottom right panel of Fig. 3, while the grey contour indicates the 95% confidence interval. It is evident that there is large scatter around the best-fit line, , where and , with a 95% confidence interval of 0.8374 to 1.1227 and -0.6859 to -0.0168 for and respectively. We tested whether the correlation between and is real using a t-test, the Spearman correlation coefficient (, with a corresponding p-value ), and bootstrapping. In all our tests we cannot reject the null hypothesis of a strong, positive correlation between the rise and the decay times.
@@
Figure 4 shows the distribution of the redshift for the TDEs in the main sample. Interestingly, this distribution also follows a normal distribution in log-space with equal to the mean and the standard deviation of the data respectively (K-S statistic equal to 0.0579 and ). We discuss this further in Sect. 5.1.
@@
Additionally, we use the Anderson-Darling (AD) normality test to further examine the validity of the log-normal behaviour of the timescale and redshift distributions. In all cases, we cannot reject the null hypothesis that the distributions are normal (in log-space).
@@ 4.2 Spectral classification
In this section we investigate the different spectral classes of the TDEs in our sample. For consistency, we adopt the three spectral classes defined in van Velzen et al. (2021), along with the TDE-featureless class introduced by Hammerstein et al. (2023):
@@
TDE-H, where the spectrum exhibits distinct and broad H and H emission lines
@@
TDE-H+He, where the spectrum shows both broad H and H emission lines and broad He ii emission features
@@
TDE-He, where the only distinct broad feature in the spectrum appears near the He ii emission line with no detectable Balmer emission lines
@@
TDE-featureless, where the spectrum primarily displays host absorption lines with no distinct emission features characteristic of the other three classes
@@
We note that for most objects in our sample only the classification spectrum is available, typically from the time of the flare. This makes the spectral classification uncertain, as the spectral properties of some TDEs evolve over time (e.g. Charalampopoulos et al. 2022). Several previous studies have already classified the TDEs in their samples, including van Velzen et al. (2021), Hammerstein et al. (2023), and Yao et al. (2023). In some cases, the resulting classifications are mixed. For example, AT 2019mha, AT 2019bhf, and AT 2018hyz appear in two or even all of the above studies, but have different classifications. Additionally, Charalampopoulos et al. (2022) provided spectral classifications for the TDEs in their sample (see their Table 4), some of which contradict the classifications assigned by van Velzen et al. (2021) for overlapping sources (namely ASASSN-15oi and PTF09ge). Furthermore, Charalampopoulos et al. (2022) identified three TDEs: AT 2018hyz, AT 2017eqx and ASASSN-14ae, that show spectral evolution over time (see their Table 4). Throughout this analysis, we assign events to the TDE-H+He class if they have been classified as such in at least one observation epoch. Additionally, 6 more transients listed in TNS have already been assigned a spectral class (3 TDE-H+He and 3 TDE-He). Moreover, spectral classifications or information on the broadness of the emission lines are available in TNS AstroNotes or classification reports of several TDEs. Finally, since the classification spectra are publicly available for almost all sources in the catalogue, we classify the remaining TDEs in this work.
@@
The results of this analysis are presented in Fig. 5. To summarize, 17.6% (23) of the TDEs belong to the TDE-H class, 12.2% (16) are classified as TDE-He, 60.3% (79) fall into the TDE-H+He class and 9.9% (13) are TDE-featureless. We note that even when considering TDEs with multiple classifications based on spectra from different epochs, the overall distribution of spectral classes remains largely unchanged. If we consider the alternative classifications (PTF09geTDE-He, ASASSN-14aeTDE-H, ASASSN-15oiTDE-He, AT 2017eqxTDE-He, AT 2018hyzTDE-H, AT2019bhfTDE-H, AT2019mhaTDE-H), the aforementioned percentages change to 20.6% (27), 13.7% (18), 55.7% (73) and 9.9% (13) for the TDE-H, TDE-He, TDE-H+He and TDE-featureless classes. The spectral classification of each TDE, along with the corresponding references, can be found on the catalogue GitHub page. We discuss the implications of our results in Sect. 5.2
@@ 4.3 Repeating flares
A few TDEs in our catalogue exhibit repeating flares. These flares may correspond to separate disruption events, where different stars are disrupted each time, or they may result from the partial disruption of the same star. In some cases, the re-brightening could be caused by emission from an accretion disc, toward the later stages of the flare. Additionally, AT 2021mhg represents a unique case believed to be a TDE followed by a supernova (SN; Somalwar et al. 2023b). Below, we present our sample of repeating-flare TDEs and our analysis of four previously unreported cases: AT 2024pvu, AT 2022exr, AT 2021uvz, and AT 2019teq. To the best of our knowledge, these TDEs have not yet been reported as exhibiting a repeating flare. We also study AT 2020acka and AT 2019ehz, who have been recently referenced briefly for having re-brightening bumps (Guo et al. 2025; Zhong 2025).
In total we identify 12 TDEs with multiple flares. The corresponding light curves highlighting the repeating outbursts are shown in Fig. 6. AT 2018fyk is plotted separately in Fig. 11, since it is the only TDE with Gaia photometry, which cannot be converted to flux. We select the best sampled light curves from the available surveys to maximize the visibility of the flares, plotting only the relevant flare regions.

### v2

@@ 3.3 Ultraviolet photometry
To evaluate TDE variability in UV-optical bands, we used data from the Swift Ultraviolet and Optical Telescope (UVOT; Roming et al. 2005), which provides photometry in three near-UV (UVW2, UVM2 and UVW1) and three optical (U, B, V) bands.
The UVOT data were downloaded from HEASARC111111https://heasarc.gsfc.nasa.gov/ and reduced using standard procedures121212http://www.swift.ac.uk/analysis/uvot/image.php (HEAsoft package v. 6.33.2). After checking the correct World Coordinate System alignment with USNO-B Catalogue (Monet et al. 2003), we combined image extensions using uvotimsum and merged exposure maps.
Sources were detected using uvotdetect. As with XRT data (see Appendix B.1), we assigned UVOT counterparts based on proximity to the TDE source coordinates: if a UVOT sources was detected within , its coordinates were adopted. If no source was detected within , we used the TDE source coordinates.
We performed source photometry using uvotsource with an aperture radius of for all filters. The background region was an annulus centred at the source with inner and outer radii of and , respectively, removing emission from overlapping sources. All magnitudes were corrected for Galactic extinction using reddening estimates from Schlafly and Finkbeiner (2011) and the extinction model from Fitzpatrick and Massa (2007). We note that these data are not host subtracted.
@@ 3.4 Spectroscopic data
All optical spectra included in the catalogue were obtained from either TNS or the Weizmann Interactive Supernova Data Repository131313https://www.wiserep.org (WISeREP; Yaron and Gal-Yam 2012). Hence, the catalogue does not include optical spectra for TDEs not classified in TNS. For the 12 TDEs from Hammerstein et al. (2023), the classification spectra can be found in their Figs. 2 and 14-16. For AT 2018lni, AT 2019cho, AT 2019mha and AT 2019meg, the classification spectra can be found in van Velzen et al. (2021).
Additionally, in the following we list references for (mainly optical) spectra of 13 TDEs that are not in the TNS-TDE sample or included in a previously published catalogue:
AT 2023vto: Fig. 4 in Kumar et al. (2024).
AT 2022agi: Figs. 2 (optical), 3 (UV) in Sun et al. (2024).
AT 2017eqx: Figs. 1, 3 in Nicholl et al. (2019).
LSQ12dyw.†141414The † symbol indicates the sources that have spectra available in WISeREP.
PTF09djl†: Figs. 4, 14 in Arcavi et al. (2014).
PTF09ge†: Fig. 12, 14 in Arcavi et al. (2014).
PS1-10jh†: Fig. 1 in Gezari et al. (2012).
iPTF15af†: Figs. 3 (optical), 4 (UV) in Blagorodnova et al. (2019).
iPTF16axa†: Figs. 7, 8 in Hung et al. (2017).
ASASSN-14li†: Fig. 3 in Holoien et al. (2016b).
ASASSN-14ae†: Figs. 4, 5 in Holoien et al. (2014).
ASASSN-15oi†: Fig. 3 in Holoien et al. (2016a).
@@ 4 Exploring the catalogue
After having constructed our main sample, we now explore its various subcategories and interesting objects. Specifically, we present the statistics and spectral classes of the TDEs in the catalogue, along with new results on TDEs with repeating flares, delayed IR emission and X-ray outbursts.
@@ 4.1 Catalogue statistics
In this section we analyse the flaring characteristics of the TDE population, in terms of rise and decay times and the event durations. To achieve this, we applied the Bayesian Blocks (Scargle 1998; Scargle et al. 2013) algorithm, which partitions a one-dimensional flux time series into blocks, each modelled by a constant flux. To avoid overfitting, we include a user-specified penalty parameter for each additional block.
Given flux measurements and their uncertainties , we first compute cumulative weighted sums (with ) up to each index: . These cumulative sums allow for efficient computation of sums over any sub-interval via . Consequently, the maximum-likelihood estimate for the constant flux in a block spanning is .
@@
Moreover, the algorithm uses dynamic programming to determine the optimal segmentation. For each index (ranging from 1 to the total number of data points ), all possible previous boundaries (from to ) are considered. We computed
@@
penalty is a user-specified constant subtracted each time a new block is added. We chose a default value of ten for the penalty to avoid overfitting or underfitting (i.e. small or large penalty values respectively).
@@
The optimal segmentation was obtained by choosing the that maximises candidate_score followed by a backtracking step (from to ) to recover the optimal boundaries.
@@
Applying the Bayesian Blocks algorithm to the flux light curves yields blocks within which the flux is assumed to be constant. For each block, we record its mean flux value. The peak of the flare is then identified as the block with the highest mean flux. We primarily use ZTF(zg) light curves (76/103). In case ZTF data are not available or the ZTF(zg) light curve is poorly sampled, we used the light curve with the best sampling. This resulted in 11 ZTF(zr) light curves, six ATLAS light curves, five ASAS-SN light curves, one GSA(G) light curve, one PS1 light curve, one CRTS(CL) light curve, and two PTF light curves.
@@
To estimate the rise and decay times of the flare, we apply a amplitude-based threshold (amplitude = peak flux - minimum flux). Starting from the peak block, we move to lower flux levels (to lower or higher block indices) until encountering a block with a mean flux below this threshold. The first sub-threshold block on the left marks the start of the rise, while the first sub-threshold segment on the right marks the end of the decay. From these boundaries, we calculate the rise time, , and decay time, , and determine their ratio . We note that in cases where the noise of the data is larger than the amplitude-based threshold, the flare is not constrained to when it drops to , but where it reaches the observational noise.
@@
An example of this process is displayed in Fig. 1. For this example, the rise time is 25698 days, the decay time 582192 days and the peak time 5980322 modified Julian date. The duration of the flare, calculated as the sum of the rise and decay times, is roughly 838 days.
@@
We applied this method to 103 TDE flares in our sample, excluding most of the sources detected in 2024 and some older ones where rise or decay times could not be reliably estimated due to large gaps in the light curves. We also included both flares for AT 2022dbl and AT 2020vdq (see Sect. 4.3). Figure 2 shows the distribution of the common logarithm (i.e. logarithm with base 10) of the ratio in our sample. A Kolmogorov-Smirnov (KS) test of the data against a normal distribution with and (the mean and standard deviation of the data respectively) yields a p-value of 0.99 (KS statistic = 0.04). This means that the ratio distribution is consistent with a log-normal distribution.
Furthermore, we estimated the statistical errors of the histogram in Fig. 2 (and Fig. 3). First, we set a fixed width for the bins of the histogram. Then we computed realisations of mock histograms by generating new values from a normal distribution of equal to each observed value (black step histogram data) and , its corresponding error. This resulted in mock histograms, with different bin heights. We computed the median of all the realisations for each bin along with its uncertainty (the distance from the median to the 16th percentile for the lower error and the distance from the 84th percentile to the median for the upper error), plotted as the grey points and grey boxes in Fig. 2 (and Fig. 3).
@@
Figure 3 displays the duration, rise and decay times of the flares. We utilised KS tests against Gaussian distributions with means151515When back-transforming from log space to linear space, smearing retransformation needs to be applied (Duan 1983). of 40030 days, 1006 days and 30020 days for the duration, and (the error is the standard error of the mean) and standard deviations of 260 days, 74 days and 229 days respectively (which correspond to the mean and standard deviation of the data). The p-values are 0.87, 0.91 and 0.58 (KS statistic of 0.057, 0.054 and 0.075) for the durations, and respectively, consistent with log-normal distributions. We repeat the same process for the statistical errors of the histograms as presented in Fig. 2 for Fig. 3.
@@
The bottom-right panel of Fig. 3 shows a scatter plot of , with error bars calculated as , where is half the width of the rise or decay blocks and is half the width of the peak block. Given errors in both the independent [] and dependent [] variables, we employed the orthogonal distance regression (ODR) to fit a best-fit line while accounting for both uncertainties. The best-fit ODR line is shown as the red solid line in the bottom-right panel of Fig. 3, while the grey contour indicates the 95% confidence interval. It is evident that there is large scatter around the best-fit line, , where and , with a 95% confidence interval of 0.768 to 1.062 and -0.65 to -0.04 for and respectively. We tested whether the correlation between and is real using a t-test, the Spearman correlation coefficient (, with a corresponding p-value ), and bootstrapping. In all our tests we could not reject the null hypothesis of a strong, positive correlation between the rise and the decay times.
@@
Figure 4 shows the redshift distributions for the TDEs in the main sample (black step histogram), as well as the optical TDEs (grey step-filled histogram). Both distributions also follow normal distributions in log-space with (full sample) equal to the mean and the standard deviation of the data respectively (KS statistic equal to 0.0579 and ) and for the optical TDEs (KS statistic equal to 0.0532 and ). We discuss further in Sect. 5.1 about potential selection effects that could cause these distribution profiles, along with their use for generating mock data.
@@
Additionally, we used the Anderson-Darling (AD) normality test to further examine the validity of the log-normal behaviour of the timescale and redshift distributions. In all cases, we could not reject the null hypothesis that the distributions are normal (in log-space).
@@ 4.2 Spectral classification
In this section we investigate the different spectral classes of the TDEs in our sample. For consistency, we adopted the three spectral classes defined in van Velzen et al. (2021) as well as the TDE-featureless class introduced by Hammerstein et al. (2023):
@@
TDE-H: In this case, the spectrum exhibits distinct and broad H and H emission lines.
@@
TDE-H+He: Here, the spectrum shows both broad H and H emission lines and broad He II emission features.
@@
TDE-He: For this case, the only distinct broad feature in the spectrum appears near the He II emission line with no detectable Balmer emission lines.
@@
TDE-featureless: Here, the spectrum primarily displays host absorption lines with no distinct emission features characteristic of the other three classes.
@@
We note that for most objects in our sample only the classification spectrum is available, typically from the time of the flare. This makes the spectral classification uncertain, as the spectral properties of some TDEs evolve over time (e.g. Charalampopoulos et al. 2022). Several previous studies have already classified the TDEs in their samples, including van Velzen et al. (2021), Hammerstein et al. (2023), and Yao et al. (2023). In some cases, the resulting classifications are mixed. For example, AT 2019mha, AT 2019bhf, and AT 2018hyz appear in two or even all of the above studies, but have different classifications. Additionally, Charalampopoulos et al. (2022) provided spectral classifications for the TDEs in their sample (see their Table 4), some of which contradict the classifications assigned by van Velzen et al. (2021) for overlapping sources (namely ASASSN-15oi and PTF09ge). Furthermore, Charalampopoulos et al. (2022) identified three TDEs: AT 2018hyz, AT 2017eqx and ASASSN-14ae, that show spectral evolution over time (see their Table 4). Throughout this analysis, we assign events to the TDE-H+He class if they have been classified as such in at least one observation epoch. Additionally, six more transients listed in TNS have already been assigned a spectral class (three TDE-H+He and three TDE-He). Moreover, spectral classifications or information on the broadness of the emission lines are available in TNS AstroNotes or classification reports of several TDEs. Finally, since the classification spectra are publicly available for almost all sources in the catalogue, we classify the remaining TDEs in this work.
@@
The results of this analysis are presented in Fig. 5. To summarise, 17.6% (23) of the TDEs belong to the TDE-H class, 12.2% (16) are classified as TDE-He, 60.3% (79) fall into the TDE-H+He class and 9.9% (13) are TDE-featureless. We note that even when considering TDEs with multiple classifications based on spectra from different epochs, the overall distribution of spectral classes remains largely unchanged. If we consider the alternative classifications (PTF09geTDE-He, ASASSN-14aeTDE-H, ASASSN-15oiTDE-He, AT 2017eqxTDE-He, AT 2018hyzTDE-H, AT 2019bhfTDE-H, AT 2019mhaTDE-H), the aforementioned percentages change to 20.6% (27), 13.7% (18), 55.7% (73) and 9.9% (13) for the TDE-H, TDE-He, TDE-H+He and TDE-featureless classes. The spectral classification of each TDE, along with the corresponding references, can be found on the catalogue GitHub page. We discuss the implications of our results in Sect. 5.2
@@ 4.3 Repeating flares
A few TDEs in our catalogue exhibit repeating flares. These flares may correspond to separate disruption events, where different stars are disrupted each time, or they may result from the partial disruption of the same star. In some cases, the re-brightening could be caused by emission from an accretion disc, towards the later stages of the flare. Additionally, AT 2021mhg represents a unique case believed to be a TDE followed by a supernova (SN; Somalwar et al. 2025b). Below, we present our sample of repeating-flare TDEs and our analysis of three previously unreported cases: AT 2024pvu, AT 2022exr, and AT 2021uvz. To the best of our knowledge, these TDEs have not yet been reported as exhibiting a repeating flare.
In total we identified 11 TDEs with multiple flares. The corresponding light curves highlighting the repeating outbursts are shown in Fig. 6. AT 2018fyk is plotted separately in Fig. 10, since it is the only TDE with Gaia photometry, which cannot be converted to flux. We select the best sampled light curves from the available surveys to maximise the visibility of the flares, plotting only the relevant flare regions.